### v1

@@ Abstract.
To understand the bugs in Solidity compiler and benefit future research, in this paper, we present the first systematic study on 533 Solidity compiler bugs. We carefully examined their characteristics (including symptoms, root causes, and distribution), and their triggering test cases. Our study leads to seven bug-revealing takeaways for Solidity compiler. Moreover, to study the limitation of Solidity compiler fuzzers and bring our findings into practical scenarios, we evaluate three Solidity compiler fuzzers on our constructed benchmark. The results show that these fuzzers are inefficient in detecting Solidity compiler bugs. The inefficiency arises from their failure to consider the interesting bug-inducing features, bug-related compilation flags, and test oracles.
conference: ACM SIGSOFT International Symposium on Software Testing and Analysis; 16-20 September, 2024; Vienna, Austriacopyright: noneconference: ; ;
@@ 1. Introduction
Existing smart contracts are mostly written in the Solidity language (over 90% (sol, 2023a)) and compiled to bytecode running on an Ethereum virtual machine (EVM). The quality of the compiler is pivotal to the programming of smart contracts, their runtime performance and trustworthiness. Like other compilers (Chaliasos et al., 2022; Le et al., 2014; Holler et al., 2012; Sun et al., 2016; Yang et al., 2011; Shen et al., 2021; Chaliasos et al., 2021; Tian et al., 2023a; Wang et al., 2023a), smart contract compiler is not immune to bugs and issues that can affect performance, reliability, and security of the compiled contracts, even when the source code is flawlessly crafted (sol, 2016, 2022, 2020b). A notable reported smart contract compiler flaw was its inadvertent transformation of a well-implemented reentrancy guard at the source code level into a malfunctioning version in the bytecode. This concealed flaw remained undetected for over two years until it was exploited by malicious entities in July 2023, culminating in a financial loss of approximately \$73.5 million (vyp, 2023). Hence, understanding the nuances of bugs in smart contract compilers, particularly those impacting the accuracy of the compilation process and the reliability of the resulting smart contracts, is essential to safeguard the integrity of these contracts. Such understanding is critical in preempting and mitigating potential vulnerabilities that could compromise the security and efficacy of smart contracts in real-world applications.
@@
  • We conduct the first systematic study on 533 Solidity compiler bugs and present a classification of their symptoms and root causes.
@@ 3.1. Bug Collection
We collect Solidity compiler bugs by extracting all closed issues labeled with the “bug” tag by the Solidity compiler developers. In total, we obtain a collection of 1,210 issues on GitHub. Then, we exclude unresolved and misreported issues from the dataset as they do not reveal any actual bugs. We also exclude issues from our dataset that are duplicate reports for the same bug in the compiler. Furthermore, to effectively understand the root cause of bugs from their code patches, we remove the GitHub issues where bugs are inadvertently fixed by unrelated commits. Additionally, we exclude issues involving the collective patching of multiple bugs, as it is difficult to isolate each bug and its corresponding patch for root cause analysis in our study. In total, 533 unique bugs are selected for subsequent analysis.
@@ 3.2.1. Symptoms
Symptom 1: Crash.  Crash refers to an unexpected termination or abnormal termination of a program during its execution. It is characterized by the sudden halting of the program’s functionality, often accompanied by error messages or system notifications. Crashes can occur due to various reasons, such as software bugs, memory leaks, resource exhaustion, etc. When a crash occurs, it signifies a failure in the software’s stability and reliability. Like other compilers(Shen et al., 2021; Chaliasos et al., 2021; Yang et al., 2011), the crash symptom is the major issue among all collected bugs in Solidity compiler, occupying 373 out of 533 bugs. More granularly speaking, the crash symptom consists of two sub-symptoms: 1) Valid Program Rejection and 2) Uninformative Error Message.
@@
Valid Program Rejection.  Given a valid test program following Solidity’s grammar, Solidity compiler rejects it with an informative error message (e.g., LABEL:listing:RVP) pointing out the reason for the compilation failure. This symptom takes up 241 out of 373 crash bugs.
@@ 5.2. Benchmark Construction
To evaluate the performance of the three Solidity compiler fuzzers, we construct a benchmark using all the bugs collected by us. Each bug in the benchmark is associated with four elements: bug description, code patch, test cases, and the version of the compiler that introduced the bug. To facilitate compiler fuzzer evaluation, for each bug, we use the nearest buggy release version as the subject for fuzzing test. We collect both the test cases and the compiler version if the bug is reproducible on the buggy version. However, we only consider reproducible bugs with crash symptoms in our fuzzer evaluation benchmark since the three fuzzers are not equipped with test oracles to identify other symptoms, such as analyzing the generated bytecode or ABI, determining the expected error messages, analyzing poor performance, or detecting runtime hangs. In total, we collect 108 reproducible bugs and 36 buggy release versions.
@@ 5.3. Seed Selection
Fuzzers have their own preferences for seeds. For instance, Fuzzol prefers small test cases with less than 1KB in size, while AFL-compiler-fuzzer prefers code templates (i.e., code snippets with placeholders) and code fragments (i.e., the concrete representation of placeholders in Solidity) extracted from test cases. In the experiment, we evaluate these fuzzers in two settings. In the first setting, for each fuzzer, we used the seed collection procedure mentioned in its documentation or original paper. Specifically, for AFL-compiler-fuzzer, we collect test cases from Solidity compiler and exclude test cases that trigger bugs in the benchmark. Then, we employ comby-decomposer (com, 2023) to decompose these test cases into fine-grained templates and fragments for invoking the code composition of AFL-compiler-fuzzer. We obtain 5,812 test cases, 25,305 fragments, and 39,978 templates. On average, each test case contains 32.9 lines of code (LOC). The seed pool construction for solfuzzer can be accomplished by executing isolate_tests, which is an automatic test case collector for solfuzzer. In total, 308 test cases are collected for solfuzzer, each containing 41.0 LOC. As for Fuzzol, we follow its experimental setup and collect seeds from its public repository (fuz, 2021). In total, 1,466 test cases are collected for Fuzzol, each containing only 13.6 LOC on average. In the second setting, we use the same seeds for all the three fuzzers. Specifically, we choose 5,812 test cases for AFL-compiler-fuzzer to build the common seed pool for the three fuzzers. This seed pool includes all unit test cases for the latest version of Solidity compiler except for bug-revealing test cases in the benchmark. We consider the number of bug-revealing test cases generated from this shared seed pool as a metric to assess the bug-detection abilities of each fuzzer concerning bugs in the benchmark. The number of bug-revealing test cases for bugs in the benchmark that are generated from the same seed pool is viewed as a metric to evaluate the bug-detection ability of each fuzzer.
@@ 5.5. Fuzzing Result
In the first experimental setting, where each fuzzer utilizes its preferred seeds, AFL-compiler-fuzzer is the most successful in bug detection, uncovering eight bugs in total. Fuzzol follows closely in second place, detecting five bugs. solfuzzer does not find any bug in this particular setting. While utilizing the seeds of AFL-compiler-fuzzer, Fuzzol detects a bug that is previously undetected by the AFL-compiler-fuzzer. However, Fuzzol does not detect this particular bug using its preferred seeds. On the other hand, solfuzzer encounters difficulties in loading these seeds and ultimately reports a timeout. Note that bugs detected by Fuzzol under two experimental settings do not overlap with bugs detected by AFL-compiler-fuzzer, suggesting that these two effective fuzzers explore distinct opportunities for revealing bugs.
@@ 5.6. Enhancing Bug Detection
Furthermore, the three chosen baselines all consider crash as the test oracle. However, code generation errors, exception handling errors, and optimization errors often manifest as non-crashes or uninformative-error-messages (according to section 4.1.1). Hence, fuzzers should examine error messages, warning messages, and outputs to effectively detect Solidity compiler bugs. Moreover, Solidity compiler fuzzers can learn from partial compiler fuzzers (Ma et al., 2023) that specialize in examining special components or compilation phrases. Our findings on bug distributions among compiler components aid fuzzers in effectively allocating their power budget.
@@ 6. Discussion
Formal Verification Component.  The formal verification component is unique to the Solidity compiler and does not exist in other compilers like C or deep-learning compilers. Such unique component comes with unique bugs that may not occur in other compilers. In our study, we found that formal verification error accounts for a significant portion of all the bugs, indicating its prevalence and importance. Noteworthy, we also found that the number of formal verification error reports has increased rapidly since 2020, corresponding to the emerging blockchain hacks recently (Figure 6). The formal verification errors contribute to 31.57% of the bugs collected in the last year (Oct/16/2022 - Oct/16/2023). The percentage is more than twice the ratio (14.07%) of formal verification errors among all bugs. This indicates that security issues in smart contracts have raised more attention to the formal verification feature of the Solidity compiler. However, the formal verification component still has much room for improvement. Due to the absence of a formal verification component, bug knowledge and testing knowledge for the existing compilers does not apply to Solidity compiler. Our study offers insights and guidance for the testing of the formal verification component in the compiler (Section 4.2 and 4.3). We call for more future research on improving the quality of the formal verification component in Solidity compiler.
@@
Summary.  The disparities between Solidity compiler and other compilers hinder transferring the knowledge of conventional compiler bugs to Solidity compiler. Unique Solidity compiler elements, such as formal verification, inline assembly, and gas optimization, introduce novel characteristics to Solidity compiler bugs. Our study brings a new perspective beyond the existing ones on bugs in traditional compilers and deep learning compilers.
@@ 7. Related Work
The most relevant works to our studies are the compiler bug studies. Certain compiler bug studies investigate the whole compiler. For instance, Sun et al. (Sun et al., 2016) conducted a comprehensive study on LLVM and GCC, analyzing their bugs in several aspects including duration, priority, code patches, locations, etc. Shen et al. (Shen et al., 2021) and Du et al. (Du et al., 2021) analyzed deep learning compiler bugs and expose their important features such as symptoms, root causes, locations, etc. Other compiler bug studies focus on bugs in a compilation stage. For instance, Chaliasos et al. (Chaliasos et al., 2021) concentrated on analyzing type-related compiler bugs in four JVM-based compilers. Zhou et al. (Zhou et al., 2021) studied the optimization-related bugs inside LLVM and GCC. This paper focuses on bugs inside the whole Solidity compiler and it is different from previous works. As shown in Figure 1, Solidity compiler have different structures than traditional compilers (e.g., LLVM and GCC) and deep learning compilers. Formal verification is a special component that contains the second most bugs. Besides, developers enact special optimizations on gas in Solidity compiler. These distinctions give rise to a novel bug distribution and unique bug characteristics. Regarding bug analysis, our primary focus lies in providing practical guidelines for future bug detection and isolation within the compiler. To our best knowledge, this is the first systematic study on Solidity compiler bugs.
@@ 9. Data Availability
To facilitate future research work and help replicate the experimental results, we release the bug classification results, benchmarks, experimental data, and the instructions for replicating the evaluation on https://github.com/haoyang9804/ISSTA24-Solidity-Study.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grant No. 61932021), and Hong Kong Research Grant Council/General Research Fund (Grant No. 16205722).

### v2

@@ Abstract.
To understand the bugs in Solidity compiler and benefit future research, in this paper, we present the first systematic study on 533 Solidity compiler bugs. We carefully examined their characteristics (including symptoms, root causes, and distribution), and their triggering test cases. Our study leads to seven bug-revealing takeaways for Solidity compiler. Moreover, to study the limitations of Solidity compiler fuzzers and bring our findings into practical scenarios, we evaluate three Solidity compiler fuzzers on our constructed benchmark. The results show that these fuzzers are inefficient in detecting Solidity compiler bugs. The inefficiency arises from their failure to consider the interesting bug-inducing features, bug-related compilation flags, and test oracles.
Solidity Compiler Bug, Empirical Study, Compiler Testing
copyright: acmlicenseddoi: 10.1145/3650212.3680362journalyear: 2024isbn: 979-8-4007-0612-7/24/09conference: Proceedings of the 33rd ACM SIGSOFT International Symposium on Software Testing and Analysis; September 16–20, 2024; Vienna, Austriabooktitle: Proceedings of the 33rd ACM SIGSOFT International Symposium on Software Testing and Analysis (ISSTA ’24), September 16–20, 2024, Vienna, Austriasubmissionid: issta24main-p889-pccs: Software and its engineering Compilersccs: Software and its engineering Software defect analysis
@@ 1. Introduction
Existing smart contracts are mostly written in the Solidity language (over 90% (sol, 2023a)) and compiled to bytecode running on an Ethereum virtual machine (EVM). The quality of the compiler is pivotal to the programming of smart contracts, their runtime performance and trustworthiness. Like other compilers (Chaliasos et al., 2022; Le et al., 2014; Holler et al., 2012; Sun et al., 2016; Yang et al., 2011; Shen et al., 2021; Chaliasos et al., 2021; Tian et al., 2023a; Wang et al., 2023a), smart contract compiler is not immune to bugs that can affect performance, reliability, and security of the compiled contracts, even when the source code is flawlessly crafted (sol, 2016, 2022, 2020b). A notable reported smart contract compiler flaw was its inadvertent transformation of a well-implemented reentrancy guard at the source code level into a malfunctioning version in the bytecode. This concealed flaw remained undetected for over two years until it was exploited by malicious entities in July 2023, culminating in a financial loss of approximately \$73.5 million (vyp, 2023). Hence, understanding the nuances of bugs in smart contract compilers, particularly those impacting the accuracy of the compilation process and the reliability of the resulting smart contracts, is essential to safeguard the integrity of these contracts. Such understanding is critical in preempting and mitigating potential vulnerabilities that could compromise the security and efficacy of smart contracts in real-world applications.
@@
  • We conduct the first systematic study on Solidity compiler bugs and present a classification of their symptoms and root causes.
@@ 3.1. Bug Collection
We collect Solidity compiler bugs by extracting all closed issues tagged as “bug” by Solidity compiler developers. In total, we obtain a collection of 1,210 issues on GitHub. Then, we exclude unresolved and misreported issues from the dataset as they do not reveal any actual bugs. We also exclude issues from our dataset that are duplicate reports for the same bug in the compiler. Furthermore, to effectively understand the root cause of bugs from their code patches, we remove the GitHub issues where bugs are inadvertently fixed by unrelated commits. Additionally, we exclude issues involving the collective patching of multiple bugs, as it is difficult to isolate each bug and its corresponding patch for root cause analysis in our study. In total, 533 unique bugs are obtained for subsequent analysis.
@@ 3.2.1. Symptoms
Symptom 1: Crash.  Crash refers to an unexpected termination or abnormal termination of a program during its execution. It manifests as the sudden halting of the program’s functionality, often accompanied by error messages or system notifications. Crashes can occur due to various reasons, such as memory leaks, resource exhaustion, etc. When a crash occurs, it signifies a failure in the software’s stability and reliability. Like other compilers(Shen et al., 2021; Chaliasos et al., 2021; Yang et al., 2011), the crash symptom is the major issue among all collected bugs in Solidity compiler, occupying 373 out of 533 bugs. More granularly speaking, the crash symptom consists of two sub-symptoms: 1) Valid Program Rejection and 2) Uninformative Error Message.
@@
Valid Program Rejection.  Given a valid test program following Solidity’s grammar, Solidity compiler rejects it with an informative error message (e.g., LABEL:listing:RVP). This symptom takes up 241 out of 373 crash bugs.
@@ 5.2. Benchmark Construction
To evaluate the performance of the three Solidity compiler fuzzers, we construct a benchmark using all the bugs collected by this work. Each bug in the benchmark is associated with four elements: bug description, code patch, test cases, and the version of the compiler that introduced the bug. To facilitate compiler fuzzer evaluation, for each bug, we use the nearest buggy release version as the subject for fuzzing test. We collect both the test cases and the compiler version if the bug is reproducible on the buggy version. However, we only consider reproducible bugs with crash symptoms in our fuzzer evaluation benchmark since the three fuzzers are not equipped with test oracles to identify other symptoms, such as analyzing the generated bytecode or ABI, determining the expected error messages, analyzing poor performance, or detecting runtime hangs. In total, we collect 108 reproducible bugs and 36 buggy release versions.
@@ 5.3. Seed Selection
Fuzzers have their own preferences for seeds. For instance, Fuzzol prefers small test cases with less than 1KB in size, while AFL-compiler-fuzzer prefers code templates (i.e., code snippets with placeholders) and code fragments (i.e., the concrete representation of placeholders in Solidity) extracted from test cases. In the experiment, we evaluate these fuzzers in two settings. In the first setting, for each fuzzer, we used the seed collection procedure mentioned in its documentation or original paper. Specifically, for AFL-compiler-fuzzer, we collect test cases from Solidity compiler and exclude test cases that trigger bugs in the benchmark. Then, we employ comby-decomposer (com, 2023) to decompose these test cases into fine-grained templates and fragments for invoking the code composition of AFL-compiler-fuzzer. We obtain 5,812 test cases, 25,305 fragments, and 39,978 templates. On average, each test case contains 32.9 lines of code (LOC). The seed pool construction for solfuzzer can be accomplished by executing isolate_tests, which is an automatic test case collector for solfuzzer. In total, 308 test cases are collected for solfuzzer, each containing 41.0 LOC. As for Fuzzol, we follow its experimental setup and collect seeds from its public repository (fuz, 2021). In total, 1,466 test cases are collected for Fuzzol, each containing only 13.6 LOC on average. In the second setting, we use the same seeds for all the three fuzzers. Specifically, we choose 5,812 test cases for AFL-compiler-fuzzer to build the common seed pool for the three fuzzers. This seed pool includes all unit test cases for the latest version of Solidity compiler except for bug-revealing test cases in the benchmark. We consider the number of bug-revealing test cases generated from this shared seed pool as a metric to assess the bug-detection abilities of each fuzzer concerning bugs in the benchmark.
@@ 5.5. Fuzzing Result
In the first experimental setting, where each fuzzer utilizes its preferred seeds, AFL-compiler-fuzzer ranks first by uncovering eight bugs. Fuzzol follows closely in second place, detecting five bugs. solfuzzer does not find any bug in this particular setting. While utilizing the seeds of AFL-compiler-fuzzer, Fuzzol detects a bug that is previously undetected by the AFL-compiler-fuzzer. However, Fuzzol does not detect this particular bug using its preferred seeds. On the other hand, solfuzzer encounters difficulties in loading these seeds and ultimately reports a timeout. Note that bugs detected by Fuzzol under two experimental settings do not overlap with bugs detected by AFL-compiler-fuzzer, suggesting that these two effective fuzzers explore distinct opportunities for revealing bugs.
@@ 5.6. Enhancing Bug Detection
Furthermore, the three chosen baselines all consider crash as the test oracle. However, code generation errors, exception handling errors, and optimization errors often manifest as non-crashes or uninformative-error-messages (according to section 4.1.1). Hence, fuzzers should examine error messages, warning messages, and outputs to effectively detect Solidity compiler bugs. Moreover, Solidity compiler fuzzers can learn from partial compiler fuzzers (Ma et al., 2023; Shen et al., 2025) that specialize in examining special components or compilation phrases. Our findings on bug distributions among compiler components aid fuzzers in effectively allocating their power budget.
@@ 6. Discussion
Formal Verification Component.  The formal verification component is unique to the Solidity compiler and does not exist in other compilers like C or deep learning compilers. Such unique component comes with unique bugs that may not occur in other compilers. In our study, we found that formal verification error accounts for a significant portion of all the bugs, indicating its prevalence and importance. Noteworthy, we also found that the number of formal verification error reports has increased rapidly since 2020, corresponding to the emerging blockchain hacks recently (Figure 6). The formal verification errors contribute to 31.57% of the bugs collected in the last year (Oct/16/2022 - Oct/16/2023). The percentage is more than twice the ratio (14.07%) of formal verification errors among all bugs. This indicates that security issues in smart contracts have raised more attention to the formal verification feature of the Solidity compiler. However, the formal verification component still has much room for improvement. Due to the absence of a formal verification component, bug knowledge and testing knowledge for the existing compilers does not apply to Solidity compiler. Our study offers insights and guidance for the testing of the formal verification component in the compiler (Section 4.2 and 4.3). We call for more future research on improving the quality of the formal verification component in Solidity compiler.
@@
Summary.  The disparities between Solidity compiler and other compilers hinder transferring the knowledge of conventional compiler bugs to Solidity compiler. Our study brings a new perspective beyond the existing ones on bugs in traditional compilers and deep learning compilers.
@@ 7. Related Work
The most relevant works to our studies are the compiler bug studies. Certain compiler bug studies investigate the whole compiler. For instance, Sun et al. (Sun et al., 2016) conducted a comprehensive study on LLVM and GCC, analyzing their bugs in several aspects including duration, priority, code patches, locations, etc. Shen et al. (Shen et al., 2021) and Du et al. (Du et al., 2021) analyzed deep learning compiler bugs and expose their important features such as symptoms, root causes, locations, etc. Other compiler bug studies focus on bugs in a compilation stage. For instance, Chaliasos et al. (Chaliasos et al., 2021) concentrated on analyzing type-related compiler bugs in four JVM-based compilers. Zhou et al. (Zhou et al., 2021) studied the optimization-related bugs inside LLVM and GCC. This paper focuses on bugs inside the whole Solidity compiler and it is different from previous works. As shown in Figure 1, Solidity compiler have different structures than traditional compilers (e.g., LLVM and GCC) and deep learning compilers. These distinctions introduce a novel bug distribution and unique bug characteristics (section 6.1). This paper reveals these distinctions and further provides practical guidelines for future bug detection and isolation within the compiler.
@@ 9. Data Availability
We release the bug classification results, benchmarks, experimental data, and the instructions for replicating the evaluation on https://github.com/haoyang9804/ISSTA24-Solidity-Study.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grant No. 61932021), Hong Kong Research Grant Council/General Research Fund (Grant No. 16205722), National Natural Science Foundation of China Grant Nos. 62322208, 62232001, and CCF Young Elite Scientists Sponsorship Program (by CAST).